OpenStudy (anonymous):

Pat, Tracy and Adrian, when all working together can paint a room in 2 hours. Pat painting alone can do the entire job in 5 hours. Tracy when working alone can do it in 6. Working alone, how long will it take Adrian to do the entire job?

OpenStudy (anonymous):

i will keep quiet

hero (hero):

p = time pat can paint the room alone t = time tracy can paint the room alone a = time adrian can paint the house alone h = hours they can paint the house together Represented by the following formula \[\frac{pta}{pt + pa + ta} = h\]

hero (hero):

If p = 5, t = 6 and h = 2, then: \[\frac{(5)(6)a}{(5)(6) + 5a + 6a} = 2\]

hero (hero):

\[\frac{30a}{30 + 11a} = 2\]

hero (hero):

From here you can solve for a

OpenStudy (zarkon):

I would write it as \[\frac{1}{5}+\frac{1}{6}+\frac{1}{a}=\frac{1}{2}\] (using work rates) then solve for a

OpenStudy (anonymous):

a = 15/2 hours for both solutions.

OpenStudy (anonymous):

Thanks everyone for the help. Figured it out after receiving the formula.